tina is twice the age of shelly. the sum of their ages is 65. how old is tina
Found 2 solutions by sofiyac, 1998Baby:
Answer by sofiyac(983)

are you sure it's not 66 years old??
let shelly be x years old, then tina is 2x years old, sum means add so add their ages together
x+2x=65
3x=65 divide each side by 3
x=65/3
now to find tina's age you multiply that by 2 so 65*2/3=43.33

Answer by 1998Baby(63)

tina is around 43